how do i solve this? x-2/2 = m + n...look for any restrictions and what is a restirction....
Answer by Alan3354(69443)

how do i solve this? x-2/2 = m + n...look for any restrictions and what is a restirction....
-------------
Consider using a ? at the end of a question, not ....
------------
how do i solve this? x-2/2 = m + n
Solve for what?
x?
m?
n?